George Kyparisis owns a company that manufactures sailboats. Actual demand for​ George's sailboats during each of the past four seasons was as​ follows:

                                                                                                            Year                                  

Season

1

2

3

4

Winter

1 comma 4801,480

1 comma 2001,200

1 comma 0801,080

960960

Spring

1 comma 5601,560

1 comma 4201,420

1 comma 6201,620

1 comma 5801,580

Summer

1 comma 0401,040

2 comma 1202,120

2 comma 0002,000

1 comma 9601,960

Fall

600600

750750

690690

520520

George has forecasted that annual demand for his sailboats in year 5 will equal

6 comma 000

sailboats.  

Based on the given data and using the multiplicative seasonal​ model, the demand level for​ George's sailboats in the spring of year 5 will be

______________________?

sailboats ​(enter a whole

​number).
Answer #1

Answer: Spring of season 5 will be: 1811

Explanation:

AC AD AE AF AG AH AI AJ AK AL AM
170 Multiplicative seasonal method
171 Season sales in year 1 sales in year 2 sales in year 3 sales in year 4 Seasonal factor Year 1 Seasonal factor Year 2 Seasonal factor Year 3 Seasonal factor Year 4 Average seasonal factor sales in year 5
172 sales in the season/yearly average sales in the season/yearly average sales in the season/yearly average sales in the season/yearly average Average of all seasonal factor of all years Average seasonal factor*1500
173 Winter 1480 1200 1080 960 1.26 0.87 0.80 0.76 0.93 1390
174 spring 1560 1420 1620 1580 1.33 1.03 1.20 1.26 1.21 1811
175 summer 1040 2120 2000 1960 0.89 1.54 1.48 1.56 1.37 2055
176 fall 600 750 690 520 0.51 0.55 0.51 0.41 0.50 745
177 Average 1170 1372.5 1347.5 1255

Excel formulas

AC AD AE AF AG AH AI AJ AK AL AM
170 Multiplicative seasonal method
171 Season sales in year 1 sales in year 2 sales in year 3 sales in year 4 Seasonal factor Year 1 Seasonal factor Year 2 Seasonal factor Year 3 Seasonal factor Year 4 Average seasonal factor sales in year 5
172 sales in the season/yearly average sales in the season/yearly average sales in the season/yearly average sales in the season/yearly average Average of all seasonal factor of all years Average seasonal factor*1500
173 Winter 1480 1200 1080 960 =AD173/AD$177 =AE173/AE$177 =AF173/AF$177 =AG173/AG$177 =AVERAGE(AH173:AK173) =AL173*1500
174 spring 1560 1420 1620 1580 =AD174/AD$177 =AE174/AE$177 =AF174/AF$177 =AG174/AG$177 =AVERAGE(AH174:AK174) =AL174*1500
175 summer 1040 2120 2000 1960 =AD175/AD$177 =AE175/AE$177 =AF175/AF$177 =AG175/AG$177 =AVERAGE(AH175:AK175) =AL175*1500
176 fall 600 750 690 520 =AD176/AD$177 =AE176/AE$177 =AF176/AF$177 =AG176/AG$177 =AVERAGE(AH176:AK176) =AL176*1500
177 Average =AVERAGE(AD173:AD176) =AVERAGE(AE173:AE176) =AVERAGE(AF173:AF176) =AVERAGE(AG173:AG176)